# Comprehensive Analysis of Relative Pressure Estimation Methods Utilizing 4D Flow MRI

### v1

@@ 2.5 Patient Cases
The final analysis utilized eight patient cases (Fig. 3ab) to assess the clinical feasibility of pressure estimation methods and to qualitatively evaluate pressure trends against what is known in the literature, including our knowledge of FL growth rate [33]. 4D flow was acquired on 3T MRI scanners (MR750, GE Medical Systems, Milwaukee, WI, USA; Ingenia, Philips, Best, Netherlands) with the following scan parameters: voxel size = ${\sim}1.5\text{ mm}\times 1.5\text{ mm}\times 2.5\text{ mm}$, average temporal resolution = 47 ms, VENC = 2.0 m/s. Agreement between the pressure estimators was also evaluated. 4D flow was acquired from a previous IRB-approved study [34] (HUM00120679). The data was segmented by thresholding patient MRA data and subsequently interpolating the resulting segmentation onto the 4D flow image space (Fig. 3cd). Inlet and outlet planes were generated in a 4D flow MATLAB toolbox [35]. FL GR was measured clinically as the change in maximum aortic diameter. Each pressure estimator was run on all eight in vivo cases. The pressure fields calculated by the fully spatial pressure estimators were analyzed for their coherence. The 4D pressure fields were also converted to pressure drops (inlet–TL outlet and inlet–FL outlet) to allow for comparisons with vWERP.
Figure 3: Two representative patient cases. (a) High-growth patient case MRA with 3D segmentation overlaid. The TL segmentation is blue, and the FL segmentation is gold. (b) Low-growth patient case MRA with 3D segmentation overlaid. (c) High-growth patient case 4D flow magnitude with velocity streamlines at peak systole overlaid. (d) Low-growth patient case 4D flow magnitude with velocity streamlines at peak systole overlaid.
@@ 3.1.1 Spatiotemporal Analysis
Although the spatiotemporal analysis was performed on nine unique spatiotemporal resolutions, only the 1.5 mm $\times$ 40 ms results are visualized in Fig. 4. Additional correlation plots and pressure drop visualizations can be found in section 1 of the supplement. Fig. 4abc presents the CFD, PPE, and STE pressure fields, respectively, at peak systole. Relative error values with no added noise are recorded in Table 1. Fig. 4f shows the correlation between the pressure fields calculated by STE, PPE, versus the CFD ground truth. The estimated pressure fields were highly consistent — both $p_{\text{STE}}$ and $p_{\text{PPE}}$ were very strongly correlated with $p_{\text{CFD}}$. PPE consistently mildly underestimated the ground truth pressure compared to STE.
Fig. 4e (top) shows the pressure drop from the aortic inlet to the TL outlet as calculated by vWERP, STE, and PPE. Table 2 summarizes the pressure drop linear regression results for the spatiotemporal analysis. All correlation coefficients were near unity for the 1.5 mm and 2.0 mm images but degraded dramatically for the 3.0 mm images. PPE correlation coefficients decreased the most dramatically for the 3.0 mm cases. Additional press drop linear regression plots can be found in subsection 1.5 of the supplement.
Figure 4: Overview of the in silico 1.5 mm $\times$ 40 ms results. (a) CFD ground truth pressure field at peak systole. (b) STE pressure field at peak systole. Aortic inlet (i) and true lumen outlet (TL4) are marked with arrow for subfigure d. (c) PPE pressure field at peak systole (d) Pressure drop from aortic inlet to TL outlet. (e) Linear regression of same pressure drop. (f) Linear regressions performed on STE and PPE pressure fields across all timesteps. The heatmaps indicate the distribution of pressure values in each 2D bin. The number of points in each 2D bin is normalized to the number of points in the entire column. The 1D histogram above the heatmap correlation plots illustrates the number of points present within each column.
@@ 3.1.2 Sensitivity Analysis Results
The sensitivity analysis for the 1.5 mm $\times$ 40 ms resolution is summarized in Fig. 5. The relative error over the cardiac cycle (11) is shown in Fig. 5a for STE and PPE. Although both methods had similar maximum errors, STE was slightly more accurate on average. Correlation heatmaps for the fully spatial analysis and the plane-to-plane pressure drop analysis at SNR = 10 are shown in Fig. 5b and d, respectively. A pressure error map for an SNR = 10 1.5 mm $\times$ 40 ms case is shown in Fig. 5c. Errors were lowest in the distal true lumen. In general, errors grew when moving proximally from the true lumen outlet. The exception to this trend was in the extremely narrow region of the proximal true lumen, which is shown in more detail in Fig. 5c (right). The range of total relative errors (12) is shown in Table 3. Averaged total relative errors can be found in subsection 1.3 of the supplement. Relative errors for STE were lower than for PPE; however, STE generally had larger ranges of errors, indicating that it was more sensitive to noise.
@@
In general, all methods were resilient to noise; only the 3.0 mm pressure traces in supplement subsection 1.6 had a noticeable spread of pressure values due to corrupted velocity fields. Furthermore, the calculated deviation values indicated that the amount of deviation from the baseline (no noise) cases increased substantially for the 3.0 mm cases. Across all resolutions and methods, the deviation values for the 1.5 mm and 2.0 mm cases remained below 4% in almost all cases and frequently below 2%. Of the 3.0 mm cases, the vWERP estimations showed the highest deviations from the baseline case, followed by STE, and then PPE a distant third place.
@@ 3.2 Flow Phantom Analysis
Figure 6: Overview of in vitro flow phantom analysis. (a) TBADENT velocity field at peak systole. (b) TBADENT STE pressure field at peak systole. (c) TBADENT PPE pressure field at peak systole. (d) Linear regression across all models and planes with no corruption. (e) Example pressure trace over the cardiac cycle. The pressure drop was calculated from the model inlet to the model outlet (i to o in Fig. 2). The “x” markers indicate where 4D flow timeframes were present. (f) Correlation heatmaps for each method. Linear regressions were run over all 50 $\Delta$SNR=10 corruption realizations per TBAD tear variation for a total of 150 unique datasets per correlation heatmap. The heatmaps indicate the distribution of pressure values in each 2D bin. The number of points in each 2D bin is normalized to the number of points in the entire column. The 1D histogram above the heatmap correlation plots illustrates the number of points present within each column.
@@ 3.2.1 Validation of Pressure Estimators
A broad overview of the flow phantom analysis results is given in Fig. 6, with TBADENT chosen as the representative TBAD tear variation. More extensive results, including those for TBADOR and TBADENT, can be found in the supplement. Qualitatively, the pressure fields estimated by STE and PPE demonstrated the features one would expect from observing the TBADENT velocity field (Fig. 6a). For instance, the high pressure impingement zone immediately following the entry tear in Fig. 6bc closely resembles the impingement zone in the corresponding FSI model of TBADENT, shown in Fig. 4c of Zimmermann et al. [20]. Both the STE (Fig. 6b) and PPE (Fig. 6c) pressure fields exhibited a large pressure drop across the FL entry tear, an increase in pressure where the FL entry tear jet impinged on the wall of the FL, and higher pressure drops from the aortic inlet (i) to each TL measurement point (TL 1–5) compared to each FL measurement point (FL 1–5). The TBADOR and TBADEXT pressure fields were similarly coherent with our expectations. Fig. 6df includes data from all TBAD tear variations. Across all TBAD tear variations, vWERP, STE, and PPE were strongly correlated ($r^{2}=0.91$) with catheter measurements (Fig. 6d), with a tendency to underestimate peak pressures (Fig. 6e). This trend holds across all pressure traces, which can be explored in further detail in subsection 2.3 of the supplement.
@@ 3.2.2 Corruption Analysis
Fig. 6f illustrates the effect of the $\Delta$SNR = 10 corruption across all 50 noise realizations for each pressure estimation method. Corruption of the velocity fields mildly degraded the average accuracy, while the correlations were moderately weaker ($r^{2}=0.69$). The pressure drop mean deviation (13) and max deviation (14) results for TBADENT are summarized in Table 4. For the FL and TL rows, deviations were averaged across all numbered pressure drops for that lumen (e.g., aortic inlet to TL1, aortic inlet to TL2, etc.). Additional tables for TBADOR and TBADEXT can be found in the supplement in subsection 2.2, and a more in-depth analysis for each pressure trace and the range of estimations can be found in the supplement subsection 2.3. As expected, the average deviation increased with noise. Average deviations were also higher in the true lumen across all models. PPE was significantly more resilient to corruption compared to vWERP and STE — the average PPE deviations were roughly half of the other method’s deviations.
Table 4: In Vitro Average Deviations Across TBADENT
@@ 3.3 Patient Cases
For all eight patient cases, PPE and STE produced pressure fields that were coherent with our expectations — relative pressure was highest in the aortic arch and steadily decreased when moving distally. Pressure drops were visibly higher in the TL, where most of the net flow was present [36]. Pressure drops in the FL were generally lower than in the TL, although the magnitude of this trend was dependent on the calculated growth rate of each patient [33]. As expected, lower FL growth rate patients had a higher pressure drop across the FL, and higher FL growth rate patients had a lower pressure drop across the FL (Fig. 7ab). The PPE, STE, and vWERP pressure drops were analyzed for agreement in Fig. 7cd, and additional comparisons can be found in section 3 of the supplement. Correlations between all pressure estimation methods were extremely strong ($r^{2}$ = 0.97–0.98). Pressure drops calculated by vWERP tended to be the highest, followed by STE ($m$ = 0.95–0.97 when compared to vWERP), and finally, PPE ($m$ = 0.87–0.93 compared to vWERP). This ordering of the results matched the outcome of the in vitro phantom analysis, where vWERP also produced the highest pressure drops.
Standard practice was to run the patient cases with nearest-neighbor upsampling by a factor of 2. At this level, pressure estimations were completed in a matter of minutes as timed by MATLAB’s timeit function. Timing results for a representative patient case were: 527.4 s for STE, 151.4 s for PPE, 192.7 s for vWERP when calculating the FL pressure drop, and 41.3 s for vWERP when calculating the TL pressure drop.
Figure 7: In vivo analysis results. (a) STE pressure field of a high-growth patient case at peak systole. (b) STE pressure field of a low-growth patient case at peak systole. (c) TL and FL correlation plot of STE vs. vWERP across all patient cases. (d) TL and FL correlation plot of PPE vs. vWERP across all patient cases.
@@ 4 Discussion
The pressure estimators vWERP, STE, and PPE were evaluated across in silico, in vitro, and in vivo TBAD datasets. The in silico analysis was performed on a CFD model with varying spatiotemporal resolution and synthetic noise levels. Notably, the in silico analysis allowed for comparisons between estimated pressure fields and the ground truth CFD solution. This allowed for a novel analysis of the spatial distributions of errors. The in vitro analysis was performed on a compliant TBAD flow phantom with three unique entry and exit tear variants. Twelve pressure catheter measurements were made in each flow phantom tear variation, combining the benefits of acquired 4D flow and significantly more ground truth pressure measurements than would be allowed clinically. The in vivo analysis was performed on eight patient cases. While no ground truth was present, this analysis allowed for a qualitative view of pressure fields and a chance to observe the relative pressure estimators in a clinical context.
@@ 4.1 Spatial Error Distribution
In the patient cases, small segments of 2–8 voxels with erroneous velocity measurements led to small regions (2–8 voxels as well) with massive pressure errors. High (but localized) erroneous velocity gradients led to corresponding localized pressure errors. In response to these findings, care was taken to pre-process patient data to remove 4D flow voxels with magnitudes/orientations significantly different from their neighbors. If STE and PPE are used clinically, care must be taken to appropriately filter out erroneous velocity measurements to ensure errors are not magnified in the corresponding relative pressure fields. Errors generally increased when moving proximally from the TL outlet, which is expected given that the relative pressure fields were shifted in reference to the TL outlet pressure.
Figure 8: Overview of in silico origin shift analysis for STE. (a) Overview of origin shift method for one voxel in 2D. Each square indicates a different shift in the origin by a subvoxel amount, forcing a different interpolation. (b) Example of the large errors in the FL and proximal TL after and unfavorable shit. (c) Relative errors over the cardiac cycle for five different origin shifts. Temporal resolution was fixed to 60 ms for each variation in spatial resolution.
@@ 4.2 Effect of Resolution
Increased spatial resolution generally increased the accuracy of the estimated pressure fields. Accuracy dipped when temporal resolution decreased mainly because the pressure estimators did not have the time frames available to capture the highly transient pressure drop at peak systole. Errors were primarily caused by underestimating pressure extrema (both systolic peaks and diastolic valleys) and a temporal shift in the latter half of the cardiac cycle. This behavior was more drastic in the 4D flow MRI acquisitions compared to the synthetic 4D flow images. The synthetic 4D flow images were constructed with instantaneous velocity measurements, which meant there was no temporal averaging. Temporal averaging, of course, occurred in the 4D flow scans. In the in silico analysis, the estimators may have missed peak pressures, but they accurately estimated the pressures based on the time frames they had access to. Meanwhile, the 4D flow scans both undersampled the peak systolic pressure drop and temporally averaged the time frames that were sampled, leading to more pronounced underestimation.

### v2

@@ 2.5 Patient Cases
The final analysis utilized eight patient cases (Fig. 3ab) to assess the clinical feasibility of pressure estimation methods and to qualitatively evaluate pressure trends against what is known in the literature, including our knowledge of FL growth rate [33]. 4D-flow was acquired on 3T MRI scanners (MR750, GE Medical Systems, Milwaukee, WI, USA; Ingenia, Philips, Best, Netherlands) with the following scan parameters: voxel size = ${\sim}1.5\text{ mm}\times 1.5\text{ mm}\times 2.5\text{ mm}$, average temporal resolution = 47 ms, VENC = 2.0 m/s. Agreement between the pressure estimators was also evaluated. 4D-flow was acquired from a previous IRB-approved study [34] (HUM00120679). The data was segmented by thresholding patient MRA data and subsequently interpolating the resulting segmentation onto the 4D-flow image space (Fig. 3cd). Inlet and outlet planes were generated in a 4D-flow MATLAB toolbox [35]. FL GR was measured clinically as the change in maximum aortic diameter. Each pressure estimator was run on all eight in vivo cases. The pressure fields calculated by the fully spatial pressure estimators were analyzed for their coherence. The 4D pressure fields were also converted to pressure drops (inlet–TL outlet and inlet–FL outlet) to allow for comparisons with vWERP.
Figure 3: Two representative patient cases. (a) High-growth patient case MRA with 3D segmentation overlaid. The TL segmentation is blue, and the FL segmentation is gold. (b) Low-growth patient case MRA with 3D segmentation overlaid. (c) High-growth patient case 4D-flow magnitude with velocity streamlines at peak systole overlaid. (d) Low-growth patient case 4D-flow magnitude with velocity streamlines at peak systole overlaid.
@@ 3.1.1 Spatiotemporal Analysis
The spatiotemporal analysis was performed on nine unique spatiotemporal resolutions. The 1.5 mm $\times$ 40 ms results are visualized in Fig. 4. Fig. 4abc presents the CFD, PPE, and STE pressure fields, respectively, at peak systole. Relative error values with no added noise are recorded in Table 1. Fig. 4f shows the correlation between the pressure fields calculated by STE and PPE versus the CFD ground truth. The estimated pressure fields were highly consistent — both $p_{\text{STE}}$ and $p_{\text{PPE}}$ were very strongly correlated with $p_{\text{CFD}}$. PPE consistently slightly underestimated the ground truth pressure compared to STE.
Fig. 4d shows the pressure drop from the aortic inlet to the TL outlet as calculated by vWERP, STE, and PPE. Table 2 summarizes the pressure drop linear regression results for the spatiotemporal analysis. All correlation coefficients were near unity for the 1.5 mm and 2.0 mm images, but degraded substantially for the 3.0 mm images. PPE correlation coefficients decreased the most for the 3.0 mm cases.
Figure 4: Overview of the in silico 1.5 mm $\times$ 40 ms results. (a) CFD ground truth pressure field at peak systole. (b) STE pressure field at peak systole. Aortic inlet (i) and true lumen outlet (TL4) are marked with arrows for subfigure d. (c) PPE pressure field at peak systole (d) Pressure drop from aortic inlet to TL outlet. (e) Linear regression of the same pressure drop. (f) Linear regressions performed on STE and PPE pressure fields across all timesteps. The heatmaps indicate the distribution of pressure values in each 2D bin. The number of points in each 2D bin is normalized to the number of points in the entire column. The 1D histogram above the heatmap correlation plots illustrates the number of points present within each column.
@@ 3.1.2 Sensitivity Analysis Results
The sensitivity analysis for the 1.5 mm $\times$ 40 ms resolution is summarized in Fig. 5. The relative error over the cardiac cycle (11) is shown in Fig. 5a for STE and PPE. Although both methods had similar maximum errors, STE was slightly more accurate on average. Correlation heatmaps for the fully spatial analysis and the plane-to-plane pressure drop analysis at SNR = 10 are shown in Fig. 5b and d, respectively. A pressure error map for an SNR = 10 1.5 mm $\times$ 40 ms case is shown in Fig. 5c. Errors were lowest in the distal true lumen. In general, errors grew when moving proximally from the true lumen outlet. The exception to this trend was in the extremely narrow region of the proximal true lumen, which is shown in more detail in Fig. 5c (right). The range of total relative errors (12) is shown in Table 3. Averaged total relative errors stayed roughly the same as noise was increased; the maximum change in average total relative error was 1.30% for PPE 2.0 mm $\times$ 60 ms when increasing from SNR=inf to SNR=10. Relative errors for STE were lower than for PPE for the 1.5 mm and 3.0 mm resolutions; however, STE generally had larger error ranges, indicating that it was more sensitive to noise.
@@
In general, all methods were resilient to noise; only the 3.0 mm pressure traces had a noticeable spread of pressure values due to noisy velocity fields. Furthermore, the calculated deviation values indicated that the spread from the baseline (no noise) cases increased substantially for the 3.0 mm cases. Across all resolutions and methods, the deviation values for the 1.5 mm and 2.0 mm cases remained below 4% in almost all cases and frequently below 2%. Of the 3.0 mm cases, the vWERP estimations showed the highest deviations from the baseline case, followed by STE, and then PPE a distant third place.
@@ 3.2 Flow Phantom Analysis
Figure 6: Overview of in vitro flow phantom analysis. (a) In vitro model STE pressure fields at peak systole. PPE pressure fields are not shown, but were nearly identical. (b) Linear regression across all models and planes with no corruption. (c) Example pressure trace over the cardiac cycle. The pressure drop was calculated from the model inlet to the model outlet (i to o in Fig. 2). The “x” markers indicate where 4D-flow timeframes were present. (d) Correlation heatmaps for each method. Linear regressions were run over all 50 $\Delta$SNR=30 corruption realizations per TBAD tear variation for a total of 150 unique datasets per correlation heatmap. The heatmaps indicate the distribution of pressure values in each 2D bin. The number of points in each 2D bin is normalized to the number of points in the entire column. The 1D histogram above the heatmap correlation plots illustrates the number of points present within each column. (e) Same analysis as in d, but with $\Delta$SNR=10
@@ 3.2.1 Validation of Pressure Estimators
A broad overview of the flow phantom analysis results is given in Fig. 6. Qualitatively, the pressure fields estimated by STE and PPE demonstrated the expected features relative to each model variation’s geometries and resulting velocity fields (Fig. 2). For instance, the high pressure impingement zone immediately following the TBADENT entry tear in Fig. 6a closely resembles the impingement zone in the corresponding FSI model of TBADENT, shown in Fig. 4c of Zimmermann et al[20]. Both the STE and PPE pressure fields exhibited a large pressure drop across the FL entry tear and higher pressure drops from the aortic inlet (i) to each TL measurement point (TL 1–5) compared to each FL measurement point (FL 1–5). The regression analysis in Fig. 6b indicated that vWERP was the most accurate pressure estimator when no noise was added. A representative pressure drop is illustrated in Fig. 6c to demonstrate what this trend in an individual pressure drop. Across all TBAD tear variations, vWERP, STE, and PPE were strongly correlated ($r^{2}=0.91$) with catheter measurements, with a tendency to underestimate peak pressures. This trend held across all pressure traces.
@@ 3.2.2 Corruption Analysis
Fig. 6d and e illustrate the effect of the $\Delta$SNR = 30 and $\Delta$SNR = 10 corruptions across all 50 noise realizations for each pressure estimation method. Corruption of the velocity fields mildly degraded the average accuracy, while the correlations were moderately weaker ($r^{2}$ = 0.69–0.70). The pressure drop mean deviation (13) and max deviation (14) results for all models are summarized in Table 4. For the FL and TL rows, deviations were averaged across all numbered pressure drops for that lumen (e.g., aortic inlet to TL1, aortic inlet to TL2, etc.). As expected, the average deviation increased with noise. Average deviations were also higher in the true lumen across all models. PPE was significantly more resilient to corruption compared to vWERP and STE — the average PPE deviations were roughly half of the other methods’ deviations.
Table 4: In Vitro Average Deviations Across All Models
@@ 3.3 Patient Cases
For all eight patient cases, PPE and STE produced pressure fields that were consistent with our expectations — relative pressure was highest in the aortic root and decreased when moving distally. Pressure drops were visibly higher in the TL, where a majority of the net flow was present [36]. Pressure drops in the FL were generally lower than in the TL, although the magnitude of this trend was dependent on the calculated growth rate of each patient [33]. As expected, lower FL growth rate patients had a higher pressure drop across the FL, and higher FL growth rate patients had a lower pressure drop across the FL (Fig. 7abcd). The PPE, STE, and vWERP pressure drops were analyzed for agreement in Fig. 7efg. Correlations between all pressure estimation methods were extremely strong ($r^{2}$ = 0.97–0.98). Pressure drops calculated by vWERP tended to be the highest, followed by STE ($m$ = 0.95–0.97 when compared to vWERP), and finally, PPE ($m$ = 0.87–0.93 compared to vWERP). This ordering of the results matched the outcome of the in vitro phantom analysis, where vWERP also produced the highest pressure drops.
Standard practice was to run the patient cases with nearest-neighbor upsampling by a factor of two. At this level, pressure estimations were completed in a matter of minutes as timed by MATLAB’s timeit function. Timing results for a representative patient case were: 527.4 s for STE, 151.4 s for PPE, 192.7 s for vWERP when calculating the FL pressure drop, and 41.3 s for vWERP when calculating the TL pressure drop.
Figure 7: In vivo analysis results. (a) STE pressure field of a high-growth patient case at peak systole. (b) STE pressure field of a low-growth patient case at peak systole. (c) STE pressure field of a second high-growth patient case at peak systole. (d) STE pressure field of a second low-growth patient case at peak systole. (e) TL and FL correlation plot of STE vs. vWERP across all patient cases. (f) TL and FL correlation plot of PPE vs. vWERP across all patient cases. (g) Fully spatial correlation plot of PPE vs. STE across all patient cases.
@@ 4 Discussion
The pressure estimators vWERP, STE, and PPE were evaluated across in silico, in vitro, and in vivo TBAD datasets. The in silico analysis was performed on a CFD model with varying spatiotemporal resolution and synthetic noise levels. Notably, the in silico analysis allowed for comparisons between estimated pressure fields and the ground truth CFD solution. This allowed for a novel analysis of the spatial distributions of errors. The in vitro analysis was performed on a compliant TBAD flow phantom with three unique entry and exit tear variants. Twelve pressure catheter measurements were made in each TBAD flow phantom tear variation, combining the benefits of acquired 4D-flow and significantly more ground truth pressure measurements than would be allowed clinically. The in vivo analysis was performed on eight patient cases. While no ground truth was present, this analysis allowed for a qualitative view of pressure fields and a chance to observe the relative pressure estimators in a clinical context.
@@ 4.1 Spatial Error Distribution
In the patient cases, small segments of 2–8 voxels with erroneous velocity measurements led to small regions (2–8 voxels as well) with large pressure errors. High (but localized) erroneous velocity gradients led to corresponding localized pressure errors. In response to these findings, care was taken to pre-process patient data to remove 4D-flow voxels with magnitudes/orientations significantly different from their neighbors. If STE and PPE are used clinically, care must be taken to appropriately filter out erroneous velocity measurements to ensure errors are not magnified in the corresponding relative pressure fields. Errors generally increased when moving proximally from the TL outlet, which is expected, given that the relative pressure fields were shifted in reference to the TL outlet pressure.
Figure 8: Overview of in silico origin shift analysis for STE. (a) Overview of the origin shift method for one voxel in 2D. Each square indicates a different shift in the origin by a subvoxel amount, forcing a different interpolation. (b) Example of the large errors in the FL and proximal TL after and unfavorable shit. (c) Relative errors over the cardiac cycle for five different origin shifts. Temporal resolution was fixed to 60 ms for each variation in spatial resolution.
@@ 4.2 Effect of Resolution
Increased spatial resolution generally increased the accuracy of the estimated pressure fields. Accuracy decreased when temporal resolution decreased, mainly because the pressure estimators did not have the time frames available to capture the highly transient pressure drop at peak systole. Errors were primarily caused by underestimating pressure extrema (both systolic peaks and diastolic valleys) and a temporal shift in the latter half of the cardiac cycle. This behavior was more drastic in the 4D-flow MRI acquisitions compared to the synthetic 4D-flow images. The synthetic 4D-flow images were constructed with instantaneous velocity measurements, which meant there was no temporal averaging. Temporal averaging, of course, occurred in the 4D-flow scans. In the in silico analysis, the estimators may have missed peak pressures, but they accurately estimated the pressures based on the time frames they had access to. Meanwhile, the 4D-flow scans both undersampled the peak systolic pressure drop and temporally averaged the time frames that were sampled, leading to more pronounced underestimation.